OpenStudy (anonymous):

z varies inversely with the product of x and y. when x=2 and y=4, z=0.5. what is the value of z when x=4 and y=9? please help, i've tried different things and i can't seem to get the correct answer. my choices are a: 45/2 b: 40/9 c:40 d: 1/9

OpenStudy (anonymous):

I will try thinking of it this way. Z varies INVERSELY with the product of x and y. This means, if x*y is more, then z will be less. At hand, we have x = 2 and y = 4 So, x*y = 2*4 = 8 And, z = 0.5 Now, when x = 4 and y = 9 x*y = 4*9 = 36 Now since the product is greater in the second case (36>8), the value of z should be lower than 0.5 Let's look at the options. a. 45/2 = 22.5 --> Not possible, it is greater than 0.5 b. 40/9 = 4.44 --> Again, greater than 0.5 c. 40 --> Much bigger than 0.5 d. 1/9 = 0.11 --> Could be possible, since it is lesser than 0.5 Now since the product is bigger in the second case, the value of z MUST be smaller than 0.5 So, I think the correct answer is D. 1/9 Hope this helps! :)

OpenStudy (anonymous):

ohhhhh okay. i think i get it. thank you.